# When LLMs get significantly worse: A statistical approach to detect model degradations

### v1

@@ 2.3 LLM Accuracy Evaluation and Optimization
Figure 1: Detecting accuracy degradation on Llama-3.1 8B Instruct based on empirical estimates and their uncertainties. The optimized model $\tilde{M}$ quantizes the KV cache and the attention and results in an empirical accuracy drop of $0.79\%$ over different datasets (Section 4). When naively treating the accuracy estimates as independent, the sampling error of the difference is overestimated, hindering the detection of actual accuracy differences (middle). Instead, we advocate to compare the per-sample scores. Our proposed exact one-sided McNemar’s test focuses on the degradation probability. In cases where there is no accuracy difference, the degradation probability is $0.5$ and we can exactly characterize its empirical distribution. This allows to detect small accuracy degradations as statistically significant, here with a $p$-value of 1.69e-05 (right).
In the field of machine learning, McNemar’s original test was previously used to assess which of either two models performs better on a single task (Meshbane and Morris, 1996; Rainio et al., 2024). Yang et al. (2025b) generalized McNemar’s test to be used in repeated cross-validation settings to decide which algorithms performs better. However, their test derives loose bounds on the null hypothesis and is thus overly conservative.
@@
Libraries such as the Language Model Evaluation Harness (lm-eval) (Gao et al., 2024) actually provide estimates of the standard error for each task individually and they also provide a per-sample score. However, using those standard errors as independent when comparing two models, overestimates the error of the accuracy difference, see Figure 1 and Equation 2. The Open LLM Leaderboard (Fourrier et al., 2024) provides a comprehensive set of accuracy benchmarks, that can all be run within the lm-eval package. Having accuracy estimates from multiple benchmarks provides the additional difficulty of deciding how to aggregate those numbers.
@@ 3.1 Exact One-Sided McNemar’s Test
The random variable $L_{M\tilde{M}}$ defined in Equation 3 links the two model evaluations of $M$ and $\tilde{M}$. Instead of thinking of two separate estimations, we can think of it as a joint accuracy experiment. We can define two probabilities:
@@ 3.2 Normal Approximation and Test Power
Our test treats $b+c$ as a constant and only relies on the exact binomial distribution. However, to have a data- and cost-efficient test it also matters that out of $N$ overall examples not unnecessarily many are discarded. To build this understanding, in this subsection we analyze the accuracy degradation $\delta:=\gamma-\beta={p_{\updownarrow}}(2{q_{\downarrow}}-1)$, which is directly linked to the degradation probability but also takes the flip probability into account. We will derive our insights based on an asymptotic analysis, which is a standard practice (Gretton et al., 2012),(Casella and Berger, 2002, Section 10.3.2)
@@ 3.3 Aggregation across Tasks
In Appendix 3 we include synthetic experiments generating scenarios for null hypothesis (${q_{\downarrow}}=1/2$ for all task for varying $T\in\mathcal{N}$) and alternative hypothesis (${q_{\downarrow}}>1/2$ for all or a single out of varying number of $T\in\mathcal{N}$ tasks) and demonstrate our theoretical insights and the pros and cons of the different aggregation strategies. In particular, we find that the pooled test is most powerful in scenarios where the degradation happens across tasks, whereas the max drop test is best when only a single task is affected. Fisher’s method strikes a good balance between these both extremes.
@@ 4 LLM Experiments
Our tests clearly flag all three INT4 checkpoints w4a16 with extremely small $p$-values, indicating strong evidence that the model degraded significantly. In fact, on the w4a16 variant of the 70B model our evaluation indicates a bug in vLLM. We spot-checked the checkpoint in transformers (Wolf et al., 2020), where the accuracy drop is much less severe. On the 8B model also the KV-FP8 variant gets clearly flagged by all tests. This is a particularly interesting case. Here, purely looking at the accuracy numbers without proper statistical tools would make a decision hard. On a few tasks, the model did not show degradation, while on MUSR it even improved slightly. However, our tests properly take sample size difference into account and the accuracy drops on the large datasets BBH, MATH, MMLU-pro are a strong signal of overall accuracy degradation. This is despite the pooled accuracy difference being only $0.78\%$, hence the baseline criterion $\hat{\delta}>2\%$ would not detect it. We further illustrate this case in Figure 1. On the 70B model, we find that KV-FP8 has a drop of $0.3\%$ and is still detected with statistical significance, showing the power of our approach, while here both baseline criteria fail to detect the accuracy.
@@ C.1 LLM Experiments
Table 5: Accuracy results of optimized Llama-3.1 8B Instruct, 3.3 70B Instruct, Mistral Small, and Llama-3.1B 8B (non-instruct) variants.
@@ D.2 Permutation-based criteria for non-binary scores
For continuous scores, we cannot use tests based on the binomial distribution anymore. For large datasets, we could use tests based on the asymptotic normality as discussed in Section 3.2. However, since some datasets are small, this would require additional care to check whether the asymptotic approximations are usable. We instead rely on a permutation-based tests that guarantees type-I error control for any sample size. Under the null hypothesis, both models should perform equally well and thus any test statistic we compute based on $\hat{L}_{M}(x_{1}),\ldots,\hat{L}_{M}(x_{N})]$ and $\hat{L}_{\tilde{M}}(x_{1}),\ldots,\hat{L}_{\tilde{M}}(x_{N})]$ should not be systematically larger than if we computed it on two lists where we randomly permute the $i$-th entry of the list for all $i$. We generally draw $m\in\mathbb{N}$ permutations (with replacement) and count with $b\in\mathbb{N}$ the cases where the simulated test statistic is larger than the original one. Then the $p$-value $p=\frac{b+1}{m+1}$ always correctly controls type-I errors (Phipson and Smyth, 2010).
@@ D.4 LLM Experiments
For GPT-OSS models (Agarwal et al., 2025), OpenAI released their own evaluation suite [^9] https://github.com/openai/gpt-oss/tree/main/gpt_oss/evals. We can use this to evaluate the the 20B model on GPQA (Rein et al., 2023) and AIME25101010https://huggingface.co/datasets/opencompass/AIME2025. Since the GPT-OSS models already come with their MoE modules in MX-FP4 precision by default, we could not find meaningful models that are further quantized. We therefore compare the 20B model against a rerun and against a version with FP8 KV cache. We also include a pruned variant which only has 7 experts per MoE layer gpt-oss-6.0b-specialized-all-pruned-moe-only-7-experts [^11] https://huggingface.co/AmanPriyanshu/gpt-oss-6.0b-specialized-all-pruned-moe-only-7-experts. However, this turns out to be way too aggressive. Furthermore, to have a model with more gradual degradation, we also include a run for which as ”optimization” we constrain the models context to 32,000 instead of 128,000 tokens.
The peculiarity of the GPT-OSS eval suite is that each example of the dataset is evaluated with three different levels of reasoning effort and each of them is run eight times with finite temperature. Thus, although each individual run results in a binary score, when aggregating over the 24 runs each examples is evaluated on, we get a more fine-grained score. Thus, we also need to apply the permutation-based tests to obtain the strictest $p$-values as outlined above. The results of the tests are presented in Table 13. While the rerun and FP8 KV cache are effectively lossless, the pruned model and the model that is artificially constrained to 32,000 context length show clear accuracy drops that are detected as statistically significant. Notice that for our experiments we use $m=10^{5}$ permutations and thus the smallest possible $p$-value for the pooled test and the max drop tests are $1/(10^{5}+1)\approx$ 1.00e-05. If one wishes to obtain even more sensitive $p$-values, one can further increase the number of permutations, but we do not foresee practical relevance for smaller $p$-values.

### v2

@@ 2.3 LLM Accuracy Evaluation and Optimization
Figure 1: Detecting accuracy degradation on Llama-3.1 8B Instruct based on empirical estimates and their uncertainties. The optimized model $\tilde{M}$ quantizes the KV cache and the attention and results in an empirical accuracy drop of $0.79\%$ over different datasets (Section 4). When naively treating the accuracy estimates as independent, the sampling error of the difference is overestimated (Miller, 2024), hindering the detection of actual accuracy differences (middle). Instead, we advocate to compare the per-sample scores. Our proposed exact one-sided McNemar’s test focuses on the degradation probability. In cases where there is no accuracy difference, the degradation probability is $0.5$ and we can exactly characterize its empirical distribution. This allows to detect small accuracy degradations as statistically significant, here with a $p$-value of 1.69e-05 (right).
In the field of machine learning, McNemar’s original test was previously used to assess which of either two models performs better on a single task (Meshbane and Morris, 1996; Rainio et al., 2024). Yang et al. (2025b) generalized McNemar’s test to be used in repeated cross-validation settings to decide which algorithms performs better. However, their test derives loose bounds on the null hypothesis and is thus overly conservative. Heineman et al. (2025) discussed dataset properties that best allow to reliably tell models apart.
@@
Libraries such as the Language Model Evaluation Harness (lm-eval) (Gao et al., 2024) actually provide estimates of the standard error for each task individually and they also provide a per-sample score. However, using those standard errors as independent when comparing two models, overestimates the error of the accuracy difference (Miller, 2024), see Figure 1 and Equation 2. The Open LLM Leaderboard (Fourrier et al., 2024) provides a comprehensive set of accuracy benchmarks, that can all be run within the lm-eval package. Having accuracy estimates from multiple benchmarks provides the additional difficulty of deciding how to aggregate those numbers.
@@ 3.1 Exact One-Sided McNemar’s Test
The random variable $L_{M,\tilde{M}}$ defined in Equation 3 links the two model evaluations of $M$ and $\tilde{M}$. Instead of thinking of two separate estimations, we can think of it as a joint accuracy experiment. We can define two probabilities:
@@ 3.2 Normal Approximation and Test Power
Our test treats $b+c$ as a constant and only relies on the exact binomial distribution. However, to have a data- and cost-efficient test it also matters that out of $N$ overall examples not unnecessarily many are discarded. To build this understanding, in this subsection we analyze the accuracy degradation $\delta:=\gamma-\beta={p_{\updownarrow}}(2{q_{\downarrow}}-1)$, which is directly linked to the degradation probability but also takes the flip probability into account. We will derive our insights based on an asymptotic analysis, which is a standard practice (Gretton et al., 2012),(Casella and Berger, 2002, Section 10.3.2).
@@ 3.3 Aggregation across Tasks
In Appendix 3 we include synthetic experiments generating scenarios for null hypothesis (${q_{\downarrow}}=1/2$ for all tasks for varying $T\in\mathcal{N}$) and alternative hypothesis (${q_{\downarrow}}>1/2$ for all or a single out of varying number of $T\in\mathcal{N}$ tasks) and demonstrate our theoretical insights and the pros and cons of the different aggregation strategies. In particular, we find that the pooled test is most powerful in scenarios where the degradation happens across tasks, whereas the max drop test is best when only a single task is affected. Fisher’s method strikes a good balance between these two extremes.
@@ 4 LLM Experiments
Our tests clearly flag all three INT4 checkpoints w4a16 with extremely small $p$-values, indicating strong evidence that the model degraded significantly. In fact, on the w4a16 variant of the 70B model our evaluation indicates a bug in vLLM. We spot-checked the checkpoint in transformers (Wolf et al., 2020), where the accuracy drop is much less severe. On the 8B model also the KV-FP8 variant gets clearly flagged by all tests. This is a particularly interesting case. Here, purely looking at the accuracy numbers without proper statistical tools would make a decision hard. On a few tasks, the model did not show degradation, while on MUSR it even improved slightly. However, our tests properly take sample size difference into account and the accuracy drops on the large datasets BBH, MATH, MMLU-pro are a strong signal of overall accuracy degradation. This is despite the pooled accuracy difference being only $0.78\%$, hence the baseline criterion $\hat{\delta}>2\%$ would not detect it. We further illustrate this case in Figure 1. On the 70B model, we find that KV-FP8 has a drop of $0.3\%$ and is still detected with statistical significance, showing the power of our approach, while here both baseline criteria fail to detect the degradation.
@@ C.1 LLM Experiments
Table 5: Accuracy results of optimized Llama-3.1 8B Instruct, 3.3 70B Instruct, Mistral Small, and Llama-3.1 8B (non-instruct) variants.
@@ D.2 Permutation-based criteria for non-binary scores
For continuous scores, we cannot use tests based on the binomial distribution anymore. For large datasets, we could use tests based on the asymptotic normality as discussed in Section 3.2. However, since some datasets are small, this would require additional care to check whether the asymptotic approximations are usable. We instead rely on a permutation-based test that guarantees type-I error control for any sample size. Under the null hypothesis, both models should perform equally well and thus any test statistic we compute based on $\hat{L}_{M}(x_{1}),\ldots,\hat{L}_{M}(x_{N})]$ and $\hat{L}_{\tilde{M}}(x_{1}),\ldots,\hat{L}_{\tilde{M}}(x_{N})]$ should not be systematically larger than if we computed it on two lists where we randomly permute the $i$-th entry of the list for all $i$. We generally draw $m\in\mathbb{N}$ permutations (with replacement) and count with $b\in\mathbb{N}$ the cases where the simulated test statistic is larger than the original one. Then the $p$-value $p=\frac{b+1}{m+1}$ always correctly controls type-I errors (Phipson and Smyth, 2010).
@@ D.4 LLM Experiments
For GPT-OSS models (Agarwal et al., 2025), OpenAI released their own evaluation suite [^9] https://github.com/openai/gpt-oss/tree/main/gpt_oss/evals. We can use this to evaluate the 20B model on GPQA (Rein et al., 2023) and AIME25101010https://huggingface.co/datasets/opencompass/AIME2025. Since the GPT-OSS models already come with their MoE modules in MX-FP4 precision by default, we could not find meaningful models that are further quantized. We therefore compare the 20B model against a rerun and against a version with FP8 KV cache. We also include a pruned variant which only has 7 experts per MoE layer gpt-oss-6.0b-specialized-all-pruned-moe-only-7-experts [^11] https://huggingface.co/AmanPriyanshu/gpt-oss-6.0b-specialized-all-pruned-moe-only-7-experts. However, this turns out to be way too aggressive. Furthermore, to have a model with more gradual degradation, we also include a run for which as ”optimization” we constrain the models context to 32,000 instead of 128,000 tokens.
The peculiarity of the GPT-OSS eval suite is that each example of the dataset is evaluated with three different levels of reasoning effort and each of them is run eight times with finite temperature. Thus, although each individual run results in a binary score, when aggregating over the 24 runs each example is evaluated on, we get a more fine-grained score. Thus, we also need to apply the permutation-based tests to obtain the strictest $p$-values as outlined above. The results of the tests are presented in Table 13. While the rerun and FP8 KV cache are effectively lossless, the pruned model and the model that is artificially constrained to 32,000 context length show clear accuracy drops that are detected as statistically significant. Notice that for our experiments we use $m=10^{5}$ permutations and thus the smallest possible $p$-values for the pooled test and the max drop test are $1/(10^{5}+1)\approx$ 1.00e-05. If one wishes to obtain even more sensitive $p$-values, one can further increase the number of permutations, but we do not foresee practical relevance for smaller $p$-values.